what is the perimeter of a square if the length of the diagonal = 5?
Answer by stanbon(48546)

what is the perimeter of a square if the length of the diagonal = 5?
Draw a square; draw a diagonal; label the sides as "x"; label the
diagonal as 5.
-----------
Use Pythagoras to get:
x^2 + x^2 = 5^2
2x^2 = 25
x = 5/sqrt2 or [5sqrt5]/2 or (5/2)sqrt5
--------------
Perimeter = 4*side
Perimeter = 4*(5/2)sqrt5
Perimeter = 10sqrt5
==================
Cheers,
Stan H.